### v1

# BioBERT-based Deep Learning and Merged ChemProt-DrugProt for Enhanced Biomedical Relation Extraction

Abstract
This paper presents a methodology for enhancing relation extraction from biomedical texts, focusing specifically on chemical-gene interactions. Leveraging the BioBERT model and a multi-layer fully connected network architecture, our approach integrates the ChemProt and DrugProt datasets using a novel merging strategy. Through extensive experimentation, we demonstrate significant performance improvements, particularly in CPR groups shared between the datasets. The findings underscore the importance of dataset merging in augmenting sample counts and improving model accuracy. Moreover, the study highlights the potential of automated information extraction in biomedical research and clinical practice.
Introduction
Biomedical literature serves as a vital conduit for various stakeholders within the scientific community, including biomedical researchers, clinicians, and database curators. Through articles, patents, and reports, these individuals disseminate their findings, contributing to the collective knowledge base of the field. However, the sheer volume of literature generated on a daily basis presents a significant challenge, hindering users’ ability to efficiently retrieve relevant information. Consequently, there is a pressing demand for innovative solutions to streamline information retrieval processes. Recognizing this need, Natural Language Processing (NLP) systems have emerged as promising tools to automate the extraction of pertinent information from biomedical texts. By leveraging computational algorithms, NLP systems aim to expedite the identification and extraction of key insights, thereby alleviating the manual burden placed on users. NLP, as a field of study, is dedicated to developing techniques that enable computers to understand and analyze human language in its unstructured form. Within the realm of NLP, Relation Extraction (RE) stands out as a crucial area of focus. RE involves the identification and characterization of relationships between entities mentioned within textual data. By discerning connections between various entities—such as genes, proteins, diseases, and treatments—RE facilitates the extraction of meaningful insights from biomedical literature. This capability holds immense potential for advancing biomedical research, clinical practice, and data curation efforts, ultimately driving innovation and improving outcomes within the healthcare domain.
In the landscape of biomedical literature analysis, a significant portion of existing systems is dedicated to the automatic recognition of mentions pertaining to genes, proteins, and chemicals within textual data. While these systems play a pivotal role in facilitating the identification of individual entities, a noticeable gap exists in their ability to extract and elucidate the intricate interactions between these entities. Indeed, a limited number of approaches have been developed to specifically target the extraction of interactions between genes/proteins and chemicals within textual data. Given this context, there is a clear imperative to delve deeper into the diverse relationships that exist between drugs, chemical compounds, and various biomedical entities, particularly genes and proteins. Systematic extraction of these relationships is essential for enabling comprehensive analysis and exploration of key biomedical properties across a spectrum of applications. By honing in on these interactions, researchers can unravel crucial insights into drug mechanisms, disease pathways, and therapeutic targets, paving the way for advancements in drug discovery, personalized medicine, and disease management. Efforts to enhance the extraction of such relationships hold immense promise for bolstering the efficacy and efficiency of biomedical research and clinical practice. By leveraging advanced Natural Language Processing (NLP) techniques and innovative methodologies, researchers can unlock the full potential of biomedical literature, accelerating the pace of scientific discovery and revolutionizing healthcare delivery.
This paper presents a methodology for improving relation extraction from biomedical texts, focusing on chemical-gene interactions. Leveraging deep learning approaches, specifically the BioBERT model and a fully connected network architecture, the study introduces a merging strategy to integrate the ChemProt and DrugProt datasets. Experimental results demonstrate significant performance enhancements, particularly in CPR groups shared between the datasets. The approach showcases the effectiveness of dataset merging in augmenting sample counts and improving model accuracy. The findings underscore the potential of automated information extraction in biomedical literature analysis and highlight avenues for future research in Natural Language Processing for biomedicine.
1 Related work
Many approaches have been explored for RE in the biomedical domain and this can be divided into 4 major classes: 1) rule-based, 2) machine learning-based 3) deep learning-based, and 4) contextualized language model-based approaches. Rule-based approaches utilize rules and patterns to identify relations between words. He, et al. [1] proposed a rule-based method to identify relations between chemical reactions synthesis by compiling two dictionaries: the first to identify the entities, and the second to identify the relations between entities that occur in the same sentence. Li, et al. [2] proposed a rule-based method where regular expressions were used to match drug names to a prescription list, followed by the use of co-location information to link the attributes to the drugs. Lowe, et al. [3] used the ChemicalTagger [4], an open-source tool that uses syntactic information to identify relations between entities.
Machine learning-based approaches use machine learning (ML) algorithms and statistical analysis to classify relations. Support vector machines (SVMs) dominated most of the early machine-learning-based approaches [5, 6, 7, 8]. Miller, et al. [9] used token context features, character type features, and semantic features with SVMs. Anick, et al. [10] used lists of n-grams with specific semantics as the features for SVM. Demner-Fushman, et al. [7] incorporated semantic information using concepts from the Unified Medical Language System [11] a repository of biomedical and clinical concepts. They exercised feature reduction through cross-validation. Zhu, et al. [12] focused on revising the learning algorithm by reformulating the SVM into a composite-kernel framework to achieve better performance.
Hybrid approaches combined linguistic pattern matching with ML techniques. Grouin, et al. [6] and Minard, et al. [13] trained an SVM and constructed linguistic patterns manually. They reported that there are advantages of the hybrid approach as linguistic patterns may confirm automatically-induced relations which helps adding confidence to the obtained results. While Yang, et al. [14] applied heuristic rules to generate candidate pairs of possible related entities that were fed into three ML models (SVM, random forest, and gradient boosted trees [15]) to classify the relations. The possible related entities were identified according to their distance from each other as defined by the number of sentence boundaries between the two entities, and multiple classifiers were developed to classify relations based on this distance.
Deep learning is a subfield of machine learning. Sahu, et al. [16] and Luo, et al. [17] utilized convolutional neural network (CNN) to automatically learn feature representations to reduce the need for engineered features. Sahu, et al. [16] applied CNN to build a Sentence-CNN which learns a single sentence-level representation for each relation, using discrete features. Lv, et al. [18] proposed the adoption of a CRF model and applied a deep learning model for feature optimization by the employment of autoencoder and sparsity limitation. Tang, et al. [19] used a hierarchical attention-based convolutional LSTM (ConvLSTM) model to construct a sentence as a multi-dimensional hierarchical sequence, to learn the local and global context information. Wei, et al. [20] proposed a model combining CNN and recurrent neural network (RNN) where they utilized Local context and semantic features as features [21].
Leveraging contextualized language models for RE in the biomedical domain has been gaining attention in the recent times. Alimova, et al. [22] conducted a comparison between three BERT-based models: BERT-uncased [23], BioBERT [24], and Clinical BERT [25] for extracting relations from clinical texts. Copara, et al [26] used a BERT-based method assessing five variations of the BERT language models, including a domain-specific model called ChemBERTa. Mahendran, et al. [27] utilized two general BERT and a BioBERT models to automatically detect relations between chemical compounds/drugs and genes/proteins. Zhang, et al. [28] proposed a hybrid method combining deep learning models with pattern-based rules and built a binary classifier by fine-tuning BioBERT. Mahendran, et al. [29] explored rule-based, deep learning based and BERT-based methods to identify Adverse drug events (ADEs) from clinical text. Mahendran, et al. [30] also proposed combining BERT with graph convolutional network (GCN) to extract information regarding chemicals and chemical reactions from chemical patents.
Zhou, et al. [31] utilized large language models (LLMs) by introducing LEAP, a framework using adaptive instructions and examples to find relationships in clinical data. Yoon, et al. [32] proposed a system that uses LLMs, a combination of weakly labeled data and knowledge bases to achieve better performance than standard methods. Yuan, et al. [33] created a new model, KeBioLM, that uses a well-established knowledge base called UMLS to understand biomedical text by incorporating existing medical knowledge. Tinn, et al. [34] and Delmas, et al. [35] explored training with limited data. Tinn, et al. used a medical vocabulary and specialized pre-training to create robust models for biomedical applications whereas Delmas, et al. explored different techniques including training with synthetic data generated by LLMs. Dunn, et al. [36] presented a simple sequence-to-sequence approach to joint named entity recognition (NER) and RE for complex hierarchical information by leveraging a fine-tuned LLM.
2 Method
2.1 Data
In this study, we employed two datasets, ChemProt and DrugProt, which contain annotations for chemical-gene interactions. These datasets were released as part of the BioCreative challenges  [^1] https://biocreative.bioinformatics.udel.edu/. Both datasets offer manual annotations of chemical and gene entities, along with their relationships, within the context of PubMed abstracts. Notably, DrugProt, released at a later date, exhibits a significantly larger volume of entity mentions and relations compared to ChemProt. This disparity is detailed in Table 1, which presents statistics on the number of abstracts, entities, and relations in each dataset. Both ChemProt and DrugProt categorize their chemical-gene relations into 22 distinct classes. These relations can be further aggregated into ten ChemProt Relation (CPR) groups. The definitions of these ten groups are provided in Table 2, originally sourced from ChemProt’s documentation. To streamline our analysis, we preprocessed both datasets to map all 22 relation categories to these 10 groups, as illustrated in Table LABEL:tab:mapping.
Table 1: ChemProt vs. DrugProt Datasets
Train Data Validation Data
ChemProt DrugProt ChemProt DrugProt
Number of Abstracts 1,020 3,500 612 750
Number of Entities 25,752 89,529 15,567 18,858
Number of Relations 6,437 17,274 3,558 3,765
Table 2: CPR Group-to-Relations Mapping
CPR Group Relations
CPR:1 PART_OF
CPR:2 REGULATOR, DIRECT_REGULATOR, INDIRECT_REGULATOR
CPR:3 UPREGULATOR, ACTIVATOR, INDIRECT_UPREGULATOR
CPR:4 DOWNREGULATOR, INHIBITOR, INDIRECT_DOWNREGULATOR
CPR:5 AGONIST, AGONIST_ACTIVATOR, AGONIST_INHIBITOR
CPR:6 ANTAGONIST
CPR:7 MODULATOR, MODULATOR_ACTIVATOR, MODULATOR_INHIBITOR
CPR:8 COFACTOR
CPR:9 SUBSTRATE, PRODUCT_OF, SUBSTRATE_PRODUCT_OF
CPR:10 NOT
2.2 Dataset merging methodology
Although the DrugProt and ChemProt datasets share similarities, DrugProt contains a larger number of relations compared to ChemProt. However, due to the presence of overlapping instances in both datasets, straightforward merging is not feasible. To address this challenge, we devised a merging strategy. Initially, we organized entities and relations based on the abstracts in which they appear, resulting in two distinct sets: one comprising all ChemProt abstracts and the other containing all DrugProt abstracts. During the merging process, if an abstract is exclusively present in one set, it is directly added to the merged set. However, in cases where an abstract exists in two versions across both sets, we attempt to merge the entities and relations from both versions. To ensure coherence, any disagreements are disregarded to prevent the introduction of noise into the merged set. With this merging strategy, we found no inconsistencies in terms of text and entity content. However, some relation conflicts were found: 63 relation conflicts were found in the training sets and 7 in the validation sets. For clarity, Table 3. provides the size of the merged dataset and the individual sizes of the ChemProt Relation (CPR) groups within it.
Table 3: Merged ChemProt-DrugProt Dataset
Merged Train Data Merged Validation Data
Number of Abstracts 3,824 1,184
Number of Entities 97,597 29,763
Number of Relations 20,401 6,450
CPR Group Training Data Validation Data
CPR:1 1,041 352
CPR:2 3,463 1,183
CPR:3 3,101 984
CPR:4 7,453 2,217
CPR:5 781 226
CPR:6 1,045 368
CPR:7 29 19
CPR:8 32 2
CPR:9 3,214 922
CPR:10 262 178
2.3 Relation extraction system
Our Relation Extraction (RE) model comprises two primary components: a BioBERT model and a fully connected top layer referred to as the ”top model.” BioBERT, developed by Lee et al. [24], is a contextual embedding model built upon the architecture of BERT. This model has been pre-trained using a vast corpus of biomedical texts, including PubMed abstracts and full-text articles from PMC. Notably, BioBERT offers two configurations: ’base’ and ’large.’ BioBERT Large’s pre-training was performed on a more extensive vocabulary that encompasses a broader spectrum of biomedical terms compared to BioBERT Base. The key distinctions between these configurations are outlined in the table below:
Table 4: Model Descriptions
BioBERT Base BioBERT Large
Number of Layers 12 24
Number of attention heads per layer 12 16
Number of units in hidden layer 768 1024
Vocabulary Original BERT Original BERT
+ custom 30K biomedical
The top model serves as the classification component in our RE model, and is based on a multi-layer fully connected network. The model begins by processing the embedding of the [CLS] token extracted from BioBERT. This embedded representation then passes through two hidden layers, each consisting of 1024 units, before reaching the output layer. To ensure compatibility with the training data, our output structure’s dimensions are aligned with the count of CPR groups present in the training data. Each entry within the vector of this structure represents the probability of the input being associated with the corresponding CPR group. We have developed two variants of the RE model, each built upon different configurations of BioBERT, as detailed in Table 4. Additionally, to comprehensively assess the system’s performance, we evaluated it both with and without the utilization of the aforementioned top model.
2.4 Experimental details
Each individual model underwent independent training on the designated training dataset, followed by evaluation using the validation dataset. To update the model parameters, we utilized the Adam optimizer with a weight decay of 0.01. We used a variable global learning rate scaling approach in which the learning rate increases linearly during the warmup stage, and is equal to the inverse squared-root of the step count after the warmup stage:
$lr=lr\_factor\cdot\min\left(step^{-0.5},step\cdot warm\_up^{-1.5}\right)$
In our experiments, we use $lr\_factor=0.0005$ and $warm\_up=1000$.
Throughout the training process, we implemented an early stopping mechanism with a patience threshold set to 6 steps. Subsequently, the best-performing model, as determined by the early stopping process, underwent evaluation using the test dataset. All three models underwent a training regimen of 5 epochs and were trained 5 times with varied random initializations.
For our reported results, CPR groups 7 and 8 were omitted due to their insufficient representation, which posed challenges for model generalization. Additionally, to augment the negative class (CPR:10), we introduced additional instances by selecting pairs of chemical and gene entities lacking any association in each input sentence. This approach contributed to the generation of more representative negative class instances.
2.5 Evaluation Metrics
We assessed the performance of our method using precision, recall, and $F_{1}$ score. Precision represents the ratio of correctly predicted mentions to the total set of predicted mentions for a specific entity, while recall signifies the ratio of correctly predicted mentions to the actual number of mentions. The $F_{1}$ score, calculated as the harmonic mean between precision and recall, provides a balanced measure of the model’s overall performance. To facilitate model evaluation, we randomly split our original merged training set into an 80/20 ratio, creating separate training and validation sets. The validation dataset was subsequently repurposed as the test set for final evaluation. During training, the 20% portion of the original merged training data served as the validation set for hyperparameter tuning and model optimization.
3 Results and Discussion
Table 5: Results for Different BioBERT-Based Models
BioBERT-Base BioBERT-Large BioBERT-Large
w/Top Model w/o Top Model w/ Top Model
P R F P R F P R F
CPR-1 0.7899 0.8607 0.8233 0.8773 0.8734 0.8749 0.8623 0.8871 0.8733
CPR-2 0.5574 0.4903 0.5202 0.5331 0.5085 0.5199 0.5380 0.5066 0.5213
CPR-3 0.8333 0.8677 0.8494 0.8713 0.8811 0.8760 0.8671 0.8809 0.8735
CPR-4 0.8697 0.9253 0.8965 0.9010 0.9310 0.9157 0.8929 0.9311 0.9116
CPR-5 0.8898 0.8956 0.8920 0.8802 0.9425 0.9093 0.8861 0.9319 0.9080
CPR-6 0.8834 0.9402 0.9105 0.9254 0.9533 0.9390 0.9276 0.9440 0.9357
CPR-9 0.7325 0.8279 0.7772 0.7928 0.8253 0.8067 0.8203 0.8157 0.8174
CPR-10 0.9292 0.9140 0.9215 0.9325 0.9256 0.9290 0.9319 0.9270 0.9294
Micro F1 0.8836 0.8948 0.8952
Std dev 0.0021 0.0026 0.0035
The $F_{1}$ scores obtained from the test set across five individual runs are presented in Table 5. These results consistently demonstrate a pattern: BioBERT-Large consistently outperforms BioBERT-Base in terms of $F_{1}$ scores. Additionally, the inclusion of the top model leads to an enhancement in the $F_{1}$ score. However, it is important to note that this enhancement is not statistically significant, indicating that achieving state-of-the-art results does not necessarily require the adoption of a more complex system.
We conducted a comprehensive comparison between the performance of our merged dataset and the ChemProt dataset processed by Sun et al. [37] For this evaluation, we specifically focused on CPR groups 3, 4, 9, and 10, as they are common to both datasets processed by Sun et al. The substantial increase in the size of our dataset has led to a significant improvement in performance. The results for various CPR groups are presented in Table 6, with Precision, Recall and $F_{1}$ scores averaged over five execution runs. Notably, our RE model demonstrates a marked increase in performance across all positive groups (CPR 3, 4, and 9), attributed to the augmented sample counts within these groups. By concentrating on CPR groups 3, 4, and 9, shared between the ChemProt and DrugProt datasets, we aim to underscore the effectiveness of our approach. These groups serve as a crucial benchmark for evaluating the impact of our merged dataset, representing interactions consistently captured across both datasets. Analyzing the performance enhancements in these shared groups highlights how increased sample counts resulting from the merger of ChemProt and DrugProt contribute to the overall effectiveness of our relation extraction model. This focused comparison demonstrates the tangible benefits of our methodology in accurately predicting chemical-gene interactions.
Table 6: Results on ChemProt vs Merged Datasets with BIOBert-Base Model
ChemProt ChemProt-DrugProt
P R F P R F
CPR-3 0.7372 0.7407 0.7384 0.8690 0.8785 0.8735
CPR-4 0.7951 0.8236 0.8083 0.8894 0.9263 0.9073
CPR-9 0.6026 0.6976 0.6410 0.7620 0.8520 0.8040
CPR-10 0.9544 0.9415 0.9479 0.9725 0.9575 0.9649
Micro F1 0.0.7839 0.8875
Std dev 0.0025 0.0021
By focusing on CPR groups 3, 4, and 9, which are common between the ChemProt and DrugProt datasets, we aim to highlight the effectiveness of our approach. These specific groups serve as a crucial benchmark for evaluating the impact of our merged dataset, as they represent interactions that are consistently captured across both datasets. By analyzing the performance improvements in these shared groups, we can demonstrate how the increased sample counts resulting from the merging of ChemProt and DrugProt contribute to enhancing the overall effectiveness of our relation extraction model. This targeted comparison allows us to showcase the tangible benefits of our methodology in capturing and accurately predicting chemical-gene interactions that are pertinent to both datasets, thus affirming the utility and robustness of our approach.
4 Conclusion & Future Work
In conclusion, this study presents a robust methodology for enhancing relation extraction from biomedical texts, with a focus on chemical-gene interactions. By integrating the ChemProt and DrugProt datasets using a novel merging strategy, we demonstrated significant improvements in model performance, particularly in CPR groups shared between the datasets. Leveraging the BioBERT model and a fully connected network architecture, our approach effectively captures and predicts complex relationships within biomedical literature. The findings underscore the importance of dataset merging in augmenting sample counts and improving model accuracy, highlighting the potential of automated information extraction in biomedical research and clinical practice. Moving forward, further exploration of Natural Language Processing techniques in biomedicine holds promise for advancing knowledge discovery and innovation in the field.
In future work, several avenues for research and development present themselves to further enhance the capabilities of relation extraction in biomedical literature analysis. Firstly, investigating advanced methods for entity recognition and disambiguation could improve the accuracy of identifying chemical-gene interactions. Additionally, exploring techniques for incorporating contextual information and domain-specific knowledge into the model architecture may lead to more robust and context-aware relation extraction systems. Moreover, the integration of multi-modal data sources, such as images and structured data, could provide complementary information to further enrich relation extraction results. Furthermore, extending the analysis to encompass additional types of biomedical entities and relationships beyond chemicals and genes would broaden the scope of the research and facilitate a more comprehensive understanding of biological systems. Finally, collaboration with domain experts and stakeholders to validate and refine the extracted relations could enhance the practical utility of relation extraction systems in real-world biomedical applications. Overall, future work in this area holds the potential to drive significant advancements in biomedical research, clinical decision-making, and drug discovery efforts.
Acknowledgements